# Breather interactions in the integrable discrete Manakov system and trigonometric Yang-Baxter maps

### v1

@@ Appendix A Derivation of the transmission coefficients
$\bar{{\mathbf{N}}}_{n}^{\prime\,\textrm{(up)}}(\bar{z}_{1})=\frac{1}{1+\tilde{g}_{n}}\bigg{[}{\mathbf{I}}_{2}+\tilde{g}_{n}\frac{\bar{z}_{1}^{2}+z_{1}^{-2}}{\bar{z}_{1}^{2}+\bar{z}_{1}^{-2}}\bigg{(}{\mathbf{I}}_{2}-\frac{\boldsymbol{\delta}_{1}\boldsymbol{\delta}_{1}^{\dagger}}{||\boldsymbol{\delta}_{1}||^{2}}\bigg{)}\bigg{]}\,,$ (A.1a)
$\bar{{\mathbf{N}}}_{n}^{\prime\,\textrm{(up)}}(i/z_{1})=\frac{1}{1+\tilde{g}_{n}}\bigg{[}{\mathbf{I}}_{2}+\tilde{g}_{n}\frac{{z}_{1}^{-2}+\bar{z}_{1}^{2}}{{z}_{1}^{-2}+{z}_{1}^{2}}\frac{\boldsymbol{\delta}_{1}\boldsymbol{\delta}_{1}^{\dagger}}{||\boldsymbol{\delta}_{1}||^{2}}\bigg{]}\,,$ (A.1b)
@@
${{\mathbf{N}}}_{n}^{\prime\,\textrm{(dn)}}(z_{1})=\frac{1}{1+\tilde{g}_{n+1}}\bigg{[}{\mathbf{I}}_{2}+\tilde{g}_{n+1}\frac{{z}_{1}^{2}+\bar{z}_{1}^{-2}}{{z}_{1}^{2}+{z}_{1}^{-2}}\bigg{(}{\mathbf{I}}_{2}-\frac{\boldsymbol{\gamma}_{1}\boldsymbol{\gamma}_{1}^{\dagger}}{||\boldsymbol{\gamma}_{1}||^{2}}\bigg{)}\bigg{]}\,,$ (A.1g)
${{\mathbf{N}}}_{n}^{\prime\,\textrm{(dn)}}(i/\bar{z}_{1})=\frac{1}{1+\tilde{g}_{n+1}}\bigg{[}{\mathbf{I}}_{2}+\tilde{g}_{n+1}\frac{\bar{z}_{1}^{-2}+z_{1}^{2}}{\bar{z}_{1}^{-2}+\bar{z}_{1}^{2}}\frac{\boldsymbol{\gamma}_{1}\boldsymbol{\gamma}_{1}^{\dagger}}{||\boldsymbol{\gamma}_{1}||^{2}}\bigg{]}\,,$ (A.1h)
@@
$\tilde{g}_{n}=4\frac{z_{1}^{-2n+2}\bar{z}_{1}^{2n+2}}{(z_{1}^{2}-\bar{z}_{1}^{2})^{2}}||\boldsymbol{\gamma}_{1}||^{2}||\boldsymbol{\delta}_{1}||^{2}\,.$ (A.2)
@@
$\bar{{\mathbf{N}}}_{n}^{\prime\,\textrm{(up)}}(z)$ $=$ $\frac{1}{1+\tilde{g}_{n}}\bigg{[}\bigg{(}1+\tilde{g}_{n}\frac{z^{2}+z_{1}^{-2}}{z^{2}+\bar{z}_{1}^{-2}}\bigg{)}{\mathbf{I}}_{2}+\tilde{g}_{n}\frac{(\bar{z}_{1}^{2}+z_{1}^{-2})(\bar{z}_{1}^{-2}-z_{1}^{-2})}{(z_{1}^{-2}-z^{-2})(z^{2}+\bar{z}_{1}^{-2})}\frac{\boldsymbol{\delta}_{1}\boldsymbol{\delta}_{1}^{\dagger}}{||\boldsymbol{\delta}_{1}||^{2}}\bigg{]}\,,$ (A.3a)
@@
${{\mathbf{N}}}_{n}^{\prime\,\textrm{(dn)}}(z)$ $=$ $\frac{1}{1+\tilde{g}_{n+1}}\bigg{[}\bigg{(}1+\tilde{g}_{n}\frac{\bar{z}_{1}^{2}+z^{-2}}{z_{1}^{2}+z^{-2}}\bigg{)}{\mathbf{I}}_{2}+\tilde{g}_{n}\frac{(\bar{z}_{1}^{2}-{z}_{1}^{2})(\bar{z}_{1}^{2}+z_{1}^{-2})}{(z^{2}-\bar{z}_{1}^{2})(z_{1}^{2}+z^{-2})}\frac{\boldsymbol{\gamma}_{1}\boldsymbol{\gamma}_{1}^{\dagger}}{||\boldsymbol{\gamma}_{1}||^{2}}\bigg{]}\,.$ (A.3d)
@@
${{\mathbf{N}}}_{n}^{\textrm{(dn)}}(z)$ $=$ $\frac{1}{1+\tilde{g}_{n}}\bigg{[}\bigg{(}1+\tilde{g}_{n}\frac{\bar{z}_{1}^{2}+z^{-2}}{z_{1}^{2}+z^{-2}}\bigg{)}{\mathbf{I}}_{2}+\tilde{g}_{n}\frac{(\bar{z}_{1}^{2}-{z}_{1}^{2})(\bar{z}_{1}^{2}+z_{1}^{-2})}{(z^{2}-\bar{z}_{1}^{2})(z_{1}^{2}+z^{-2})}\frac{\boldsymbol{\gamma}_{1}\boldsymbol{\gamma}_{1}^{\dagger}}{||\boldsymbol{\gamma}_{1}||^{2}}\bigg{]}\,.$ (A.5c)
@@
$\bar{{\mathbf{c}}}(z)=\lim_{n\rightarrow-\infty}\frac{1}{1+re^{-4a_{1}n}}\bigg{[}\bigg{(}1+re^{-4a_{1}n}\frac{z^{2}+z_{1}^{-2}}{z^{2}+\bar{z}_{1}^{-2}}\bigg{)}{\mathbf{I}}_{2}+re^{-4a_{1}n}\frac{(\bar{z}_{1}^{2}+z_{1}^{-2})(\bar{z}_{1}^{-2}-z_{1}^{-2})}{(z_{1}^{-2}-z^{-2})(z^{2}+\bar{z}_{1}^{-2})}\frac{\boldsymbol{\delta}_{1}\boldsymbol{\delta}_{1}^{\dagger}}{||\boldsymbol{\delta}_{1}||^{2}}\bigg{]}\,,$ (A.6)
${\mathbf{c}}(z)=\lim_{n\rightarrow-\infty}\frac{1}{1+re^{-4a_{1}n}}\bigg{[}\bigg{(}1+re^{-4a_{1}n}\frac{\bar{z}_{1}^{2}+z^{-2}}{z_{1}^{2}+z^{-2}}\bigg{)}{\mathbf{I}}_{2}+re^{-4a_{1}n}\frac{(\bar{z}_{1}^{2}-{z}_{1}^{2})(\bar{z}_{1}^{2}+z_{1}^{-2})}{(z^{2}-\bar{z}_{1}^{2})(z_{1}^{2}+z^{-2})}\frac{\boldsymbol{\gamma}_{1}\boldsymbol{\gamma}_{1}^{\dagger}}{||\boldsymbol{\gamma}_{1}||^{2}}\bigg{]}\,.$ (A.7)
@@ Appendix B Long-time asymptotics for the 2-fundamental breather solution
$\bar{z}_{1}^{2n}z_{1}^{-2n}\bar{{\mathbf{C}}}_{1}(\tau){\mathbf{C}}_{1}(\tau)$ $=$ $e^{-2\zeta_{1}}\bar{z}_{1}^{2}||\boldsymbol{\gamma}_{1}||^{2}||\boldsymbol{\delta}_{1}||^{2}\boldsymbol{\Pi}\,,$ (B.1)
${z}_{1}^{-2n}\bar{z}_{1}^{2n}\hat{{\mathbf{C}}}_{1}(\tau)\tilde{{\mathbf{C}}}_{1}(\tau)$ $=$ $-e^{-2\zeta_{1}}z_{1}^{-2}||\boldsymbol{\gamma}_{1}||^{2}||\boldsymbol{\delta}_{1}||^{2}({\mathbf{I}}_{2}-\boldsymbol{\Pi})\,,$ (B.2)
@@
$\boldsymbol{\Pi}\equiv\frac{\boldsymbol{\delta}_{1}\boldsymbol{\delta}_{1}^{\dagger}}{||\boldsymbol{\delta}_{1}||^{2}},\;\;\;\boldsymbol{\Pi}^{2}=\boldsymbol{\Pi}\,.$ (B.3)
@@
$a=4\frac{z_{1}^{2}\bar{z}_{1}^{2}}{(\bar{z}_{1}^{2}-z_{1}^{2})^{2}}e^{-2\zeta_{1}}||\boldsymbol{\gamma}_{1}||^{2}||\boldsymbol{\delta}_{1}||^{2},\;\;\;b=c^{*}=4\frac{\bar{z}_{1}^{-2}z_{1}^{-2}}{(\bar{z}_{1}^{2}+\bar{z}_{1}^{-2})(\bar{z}_{1}^{-2}-z_{1}^{-2})}e^{-2\zeta_{1}}||\boldsymbol{\gamma}_{1}||^{2}||\boldsymbol{\delta}_{1}||^{2}\,.$ (B.6)
@@
Upon plugging in the definition of $a$ from (B.6), after simplification one can verify that (B.11) is identical to the 1-fundamental breather solution in (2.29a).
Although conceptually similar, computing the long-time asymptotics in the opposite limits is quite cumbersome because now the system has growing exponential terms, and at least 2 next-to-leading order terms have to be computed and retained at each step, since the system becomes degenerate. We give below some details on how the limit is computed for fixed $\zeta_{2}$ and $\tau\to-\infty$ [where, according to (4.9), $e^{-\zeta_{1}}\to+\infty$] in the case in which both solitons are fundamental solitons. The basic idea is to expand the coefficients of the system (4.1) with $J=2$, ${\mathbf{C}}_{j}=\boldsymbol{\gamma}_{j}\boldsymbol{\delta}_{j}^{\dagger}$ for $j=1,2$ and $\boldsymbol{\delta}_{j}=(1,0)^{T}$ in powers of $e^{-\zeta_{1}}$ [i.e. $\mathcal{O}(1),\mathcal{O}(e^{\zeta_{1}}),\dots$], iteratively solving the equations for one of the unknowns and back-substituting into the other equations, while keeping at least two non-zero terms at each step. This yields:

### v2

@@ Appendix A Derivation of the transmission coefficients
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Appendix B Long-time asymptotics for the 2-fundamental breather solution
[…]
@@
[…]
@@
[…]
@@
Upon plugging in the definition of $a$ from (B.6), after simplification one can verify that (B.11) is identical to the 1-fundamental breather solution in (2.29).
Although conceptually similar, computing the long-time asymptotics in the opposite limits is quite cumbersome because now the system has growing exponential terms, and at least 2 next-to-leading order terms have to be computed and retained at each step, since the system becomes degenerate. We give below some details on how the limit is computed for fixed $\zeta_{2}$ and $\tau\to$-$\infty$ (where, according to (4.9), $e^{-\zeta_{1}}\to+\infty$) in the case in which both solitons are fundamental solitons. The basic idea is to expand the coefficients of the system (4.1) with $J=2$, ${\mathbf{C}}_{j}=\boldsymbol{\gamma}_{j}\boldsymbol{\delta}_{j}^{\dagger}$ for $j=1,2$ and $\boldsymbol{\delta}_{j}=(1,0)^{T}$ in powers of $e^{-\zeta_{1}}$ (i.e. $\mathcal{O}(1),\mathcal{O}(e^{\zeta_{1}}),\dots$), iteratively solving the equations for one of the unknowns and back-substituting into the other equations, while keeping at least two non-zero terms at each step. This yields: